Head-First Company now sells both bicycle helmets and motorcycle helmets. Next year, Head- First expects to produce total revenue of $575,000 and incur total variable cost of $400,000. Total fixed cost is expected to be $60,500.

Required:
1. Calculate the break-even point in sales dollars for Head-First. Round the contribution margin ratio to four decimal places and sales to the nearest dollar.
2. Check your answer by preparing a contribution margin income statement
Answer #1

1. Contribution margin% = (sales-variable cost)/sales

= (575000-400000)/575000 = 30.4348%

Break even sales = Fixed expense / Sales

= 60500/30.4348%

= $198786

2.

Sales 198786
Less: variable cost (400000/575000*198786) 138286
Contribution margin 60500
Less: fixed expense 60500
Profit 0